OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)?

OpenStudy (anonymous):

options: x2 – 5x + 14 x2 + 5x – 14 x2 + 5x – 5 x2 + 9x + 5

OpenStudy (australopithecus):

well do you understand the notation f(x), g(x), s(x)...

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

You have to just multiply the two expressions..

OpenStudy (anonymous):

\[f(x) \cdot g(x) = (x+7)(x-2) = ??\]

OpenStudy (anonymous):

which method would I use to multiply the 2 expressions?

OpenStudy (anonymous):

Have you heard about FOIL ??

OpenStudy (anonymous):

oh yes, first, outer, inner last

OpenStudy (anonymous):

Then go for it..

OpenStudy (australopithecus):

f(x), g(x), s(x)... they all represent y, the out put of the function so instead of saying y = x + 7 y = x – 2 we replace y with the notation f(x), g(x), s(x), as if we wanted to talk about these functions it would be confusing to call them equation y the x in brackets is the value we inputed in the function so if we wanted to talk about a specific point in a function we can just right f(2), instead of 2-2 (note this is more handy when dealing with ugly functions) so just imagine f(x), g(x), s(x) as variables that represent the equations so if we say f(x) * g(x) we are just saying (x + 7)(x – 2) if we say f(g(x)) we are just replacing x in x + 7 with x – 2 so we get f(g(x)) = (x-2) + 7

OpenStudy (australopithecus):

hope that clears up the notation for you remember (x,f(x)) now instead of (x,y)

OpenStudy (anonymous):

What did you get after multiplying them @brittneyy18rm

OpenStudy (australopithecus):

(x + 7)(x – 2) = x(x) + x(-2) + 7(x) + 7(-2)

OpenStudy (anonymous):

after I multiplied them, I got x^2+5x-14 :)

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

thanks for your help!

OpenStudy (anonymous):

Welcome dear.